A certain orange beverage contains only 10% juice. How much juice has to be added to one litre of this drink so that the new mixture will contain 40% juice?
a) 0.2 litre b) 0.3 litre c) 0.4 litre d) 0.5 litre e) 0.6 litre

Answer by stanbon(75887)

A certain orange beverage contains only 10% juice. How much juice has to be added to one litre of this drink so that the new mixture will contain 40% juice?
----------
Equation:
juice + juice = juice
0.10(1) + 1*x = 0.40(1+x)
Multiply thru by 100 to get:
10 + 100x = 40 + 40x
60x = 30
x = 1/2 litre (amt. of 100% juice needed in the mixture)
======================
Cheers,
Stan H.


a) 0.2 litre b) 0.3 litre c) 0.4 litre d) 0.5 litre e) 0.6 litre